OpenStudy (anonymous):

a canon on a hill, 50m above sealevel fires a shot with v=600m/s to a ship with distance 12000m. At which angle should the shot be fired?

OpenStudy (anonymous):

i don't understand how i can know if the shot is fired up, horizontal or down...?

OpenStudy (anonymous):

First check if fired horizontal can it reach there. If it can, it is fine. Otherwise, it has to be fired at an angle. Can you try?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Are you able to proceed?

OpenStudy (anonymous):

need to find the right formulas

OpenStudy (anonymous):

if fired horizontally: x=v.t and y=y-1/2gt^2 ?

OpenStudy (anonymous):

Yeah, first find t from second equation. Then use the first. Check whether you get x=12000

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

t=3,19

OpenStudy (anonymous):

impossibru

OpenStudy (anonymous):

so i guess it's fired upwards?

OpenStudy (anonymous):

good, it should be fired upwards. I hope you can continue now. If any help,ask.

OpenStudy (anonymous):

thank you! i'll try completing it!

OpenStudy (anonymous):

Welcome, be careful with the sign convention.

OpenStudy (anonymous):

may i ask you for an answer to compare to mine?

OpenStudy (anonymous):

Okay, can you wait? Will return soon with an answer.

OpenStudy (anonymous):

ofcourse

OpenStudy (anonymous):

my answer is around 19,27 degrees

OpenStudy (anonymous):

What was the equation you got?

OpenStudy (anonymous):

i probably calculated something wrong

OpenStudy (anonymous):

i got : 12000=600.cosx.600.sinx.2/9.8

OpenStudy (anonymous):

What was your approach?

OpenStudy (anonymous):

well.. x=vi.cosx.t and y=yi + visinx - 1/2gt^2

OpenStudy (anonymous):

solve y for t gives me : vi.sinx.2/g = t

OpenStudy (anonymous):

i plug it into my x equation

OpenStudy (anonymous):

The second equation should have a t multiplied to vi. sin x .

OpenStudy (anonymous):

yes, forgot

OpenStudy (anonymous):

so i get: x=vi.cosx.visinx.2/g

OpenStudy (anonymous):

we have a quadratic in t there, am i correct?

OpenStudy (anonymous):

which t?

OpenStudy (anonymous):

here y=yi + visinx - 1/2gt^2 ?

OpenStudy (anonymous):

i factored it out so i can get vi.sinx.2/g = t

OpenStudy (anonymous):

because y = 0

OpenStudy (anonymous):

Yes, it is \[s = ut+ \frac{ 1 }{ 2 }a t^2\] you forgot the t multiplied to u

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes, but i just forgot to type it. i included it into my calculations

OpenStudy (anonymous):

what's s and u?

OpenStudy (anonymous):

am i doing something wrong?

OpenStudy (anonymous):

s- displacement u-initial velocity How did you factorize the equation?

OpenStudy (anonymous):

\[-50 = 600t \sin x -\frac{ 1 }{2 } 10 t^2\]

OpenStudy (anonymous):

i see what i did wrong here

OpenStudy (anonymous):

forgot about the 50

OpenStudy (anonymous):

and \[600\cos x \times t = 12000\]

OpenStudy (anonymous):

it's a quadratic equation

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

Substitute the t from second equation into the first. But I got stuck in solving the equation.

OpenStudy (anonymous):

so t=12000/600cosx in the y equation

OpenStudy (anonymous):

Yeah. Try solving.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

After using calculator, and a quadratic equation in sin 2x. I got 9.66 degrees. But not sure. I think its better to stop here,as we have the idea of solving but the equation is not good.

OpenStudy (anonymous):

we're using the a wrong equation?

OpenStudy (anonymous):

No, a bad equation. Looks difficult to solve. Hope there is some easier way.

OpenStudy (anonymous):

yea, i'm way off

OpenStudy (anonymous):

i'll just skip this question

OpenStudy (anonymous):

thought it would be easy

OpenStudy (anonymous):

okay. if you get a easier solution, please inform me. Solving completely is not important, the idea of how to solve matters.

OpenStudy (anonymous):

alright, thanks!

OpenStudy (anonymous):

You're welcome:)